Use modified Euler’s method with one step to find the value of y at x = 0.1 to five significant figures, where dy/dx = x^2+y, y=0.94, when x = 0.


1
Expert's answer
2021-02-02T05:21:51-0500

Solution

Modified Euler is a  method for numerical integration of ODE. If

y’(x) = f(x,y(x)), y(x0)=y0 

the size of every step h and xn = x0+n*h an approximation of the solution to the ODE is

yn+1 = yn + h*f(xn+h/2,yn+h*f(xn,yn)/2) 

For given ODE

f(x,y) = x2+y;  y(0) = 0.94;  x0 = 0;  y0 = 0.94;  h = 0.1

yn+1 = yn + h*[ yn +h*( yn+ xn2)/2+ (xn+h/2)2)]

From this expression we’ll get  x1 = 0.1;  y1 = 1.03895  

Rounded to five significant figures y1 = 1.0390

Answer

y1 = 1.0390